# Quantum dual extended Hamming code immune to collective coherent errors

### v4

@@ Abstract
Collective coherent (CC) errors are inevitable, as every physical qubit undergoes free evolution under its kinetic Hamiltonian. These errors can be more damaging than stochastic Pauli errors because they affect all qubits coherently, resulting in high-weight errors that standard quantum error-correcting (QEC) codes struggle to correct. In quantum memories and communication systems, especially when storage durations are long, CC errors often dominate over stochastic noise. Trapped-ion platforms, for example, exhibit strong CC errors with minimal stochastic Pauli components. In this work, we address the regime where immunity to CC errors, high code rate (due to limited qubit availability), and moderate distance (sufficient for correcting low-weight errors) are all essential. We construct a new family of constant-excitation (CE) stabilizer codes with parameters $2^{r+1},2^{r}-(r+1),4]]$. The smallest instance, the $[[8,1,4]]$ code, improves the code rate and error threshold of the best previously known CE code by factors of approximately two and four, respectively.
@@ I Introduction
In our previous work [8], we made significant progress in increasing the code distance without sacrificing too much code rate. The family of $[[2(w+1)(w+K),K]]$ CE codes, referred to as w-codes, can be interpreted as $[[2(w+1)(w+K),K,w+1]]$ Shor-like codes with distance $d=w+1$. Here, the parameter w indicates the ability to correct weight-w amplitude-damping (AD) errors, though not weight-w Pauli errors. For conventional $[[n,k,d]]$ stabilizer codes designed to correct Pauli errors, up to $\lfloor\frac{d}{2}\rfloor$ errors can be detected and $\lfloor\frac{d-1}{2}\rfloor$ errors can be corrected.
@@
As a result, we propose designing an outer code that does not differentiate between $X$ and $Z$ errors, relying instead on the inner CDR code to enable effective discrimination between these error types. In this work, we introduce a new family of CE stabilizer codes with parameters $2^{r+1},2^{r}-(r+1),4]]$. These codes are constructed by encoding classical extended Hamming codes $^{r},2^{r}-(r+1),4] [22$] into CE subspaces using an inner CDR code.
The smallest instance in this family is the $[[8,1,4]]$ code, which uses fewer qubits than previously known CE codes, whether stabilizer or non-stabilizer. Moreover, this new construction improves the asymptotic code rate from $\frac{1}{8}$ in our earlier work [8] to $\frac{1}{2}$. A summary comparison is provided in Table 1.
Table 1: Comparison of representative CE QEC codes. Each $[[n,k,d]]$ code encodes $k$ logical qubits into $n$ physical qubits with distance $d$, allowing correction of up to $\lfloor\frac{d-1}{2}\rfloor$ and detection of up to $\lfloor\frac{d}{2}\rfloor$ Pauli errors. The rate is defined as the ratio $\frac{k}{n}$, indicating the encoding efficiency.
@@
[[2d(K+d1),K,d]][[2d(K+d-1),K,d]] [8] $\frac{1}{2d}\frac{K}{K+d-1}$ CSS code
@@
$2^{r+1},2^{r}-(r+1),4]]$ $\frac{1}{2}-\frac{r+1}{2^{r+1}}$ Highest rate
@@ II.2 Classical extended Hamming code
$g_{0}^{Z}$ $=\bigotimes_{j=2^{r-1}}^{2^{r}-1}Z_{2^{r}-j},$ (5a)
$g_{i}^{Z}$ $=\bigotimes_{j=0}^{2^{i}-1}\bigotimes_{k=0}^{2^{r-i}-1}Z_{2^{r}-(j(2^{r-i})+k)},\quad i\in\{1,\dots,r\}.$ (5b)
@@
which can detect a single error of one type. In fact, this code is of distance $4$.
@@ III Quantum dual extended Hamming code
$g_{0}=$ $\bigotimes_{j=2^{r-1}}^{2^{r}-1}\;X_{j}\,Z_{2^{r}-j}\,X_{j+2^{r}},$ (7a)
$g_{i}=$ $\bigotimes_{j=0}^{2^{i}-1}\bigotimes_{k=0}^{2^{r-i}-1}\;X_{j(2^{r-i})+k}\,Z_{2^{r}-(j(2^{r-i})+k)}\,X_{j(2^{r-i})+k+2^{r}},$
@@
logical qubits.
@@ IV Conclusion
In this work, we considered scenarios where only a limited number of high-quality qubits are available, such as in quantum memories and communication systems subject to unavoidable CC errors and minor stochastic Pauli errors. To address these challenges, we introduced a new family of CE stabilizer codes with parameters $2^{r+1},2^{r}-(r+1),4]]$, designed to mitigate both error types while optimizing code rate. The smallest member of this family, the $[[8,1,4]]$ CE code, reduces the required qubit count by half compared to the previously known CE code and improves the error threshold by a factor of four.

### v5

@@ Abstract
Collective coherent (CC) errors are inevitable, as every physical qubit undergoes free evolution under its kinetic Hamiltonian. These errors can be more damaging than stochastic Pauli errors because they affect all qubits coherently, resulting in high-weight errors that standard quantum error-correcting (QEC) codes struggle to correct. In quantum memories and communication systems, especially when storage durations are long, CC errors often dominate over stochastic noise. Trapped-ion platforms, for example, exhibit strong CC errors with minimal stochastic Pauli components. In this work, we address the regime where immunity to CC errors, high code rate (due to limited qubit availability), and moderate distance (sufficient for correcting low-weight errors) are all essential. We construct a family of constant-excitation (CE) stabilizer codes with parameters $2^{r+1},2^{r}-(r+1),3]]$. The smallest instance, the $[[8,1,3]]$ code, improves the code rate and error threshold of the best previously known CE code by factors of approximately two and four, respectively.
@@ I Introduction
In our previous work [8], we made significant progress in increasing the code distance without sacrificing too much code rate. The family of $[[2(w+1)(w+K),K]]$ CE codes, referred to as w-codes, can be interpreted as $2(w+1)(w+K),K,d^{\prime}]]$ Shor-like codes with distance $d^{\prime}=w+1$ for $K=1$, $d^{\prime}=2$ for $K\geq 2$. Here, the parameter w indicates the ability to correct weight-w amplitude-damping (AD) errors, though not weight-w Pauli errors. For conventional $[[n,k,d]]$ stabilizer codes designed to correct Pauli errors, up to $\lfloor\frac{d}{2}\rfloor$ errors can be detected and $\lfloor\frac{d-1}{2}\rfloor$ errors can be corrected.
@@
As a result, we propose designing an outer code that does not differentiate between $X$ and $Z$ errors, relying instead on the inner CDR code to enable effective discrimination between these error types. In this work, we introduce a new family of CE stabilizer codes with parameters $2^{r+1},2^{r}-(r+1),3]]$. These codes are constructed by encoding classical extended Hamming codes $^{r},2^{r}-(r+1),4] [22$] into CE subspaces using an inner CDR code.
The smallest instance in this family is the $[[8,1,3]]$ code, which uses fewer qubits than previously known CE codes, whether stabilizer or non-stabilizer. Moreover, this new construction improves the asymptotic code rate to $\frac{1}{2}$. A summary comparison is provided in Table 1.
Table 1: Comparison of representative CE QEC codes. Each $[[n,k,d]]$ code encodes $k$ logical qubits into $n$ physical qubits with distance $d$, allowing correction of up to $\lfloor\frac{d-1}{2}\rfloor$ and detection of up to $\lfloor\frac{d}{2}\rfloor$ Pauli errors. The rate is defined as the ratio $\frac{k}{n}$, indicating the encoding efficiency. The symbol $d^{\prime}=d$ for $K=1$, otherwise $d^{\prime}=2$ for $K\geq 2$.
@@
$2d(K+d-1),K,d^{\prime}]] [$8] $\frac{1}{2d}\frac{K}{K+d-1}$ CSS code
@@
$2^{r+1},2^{r}-(r+1),3]]$ $\frac{1}{2}-\frac{r+1}{2^{r+1}}$ Highest rate
@@ II.2 Classical extended Hamming code
$g_{0}^{Z}$ $=\bigotimes_{j=2^{r-1}}^{2^{r}-1}Z_{2^{r}-1-j},$ (5a)
$g_{i}^{Z}$ $=\bigotimes_{j=0}^{2^{i}-1}\bigotimes_{k=0}^{2^{r-i}-1}Z_{2^{r}-1-(j(2^{r-i})+k)},\quad i\in\{1,\dots,r\}.$ (5b)
@@
which can detect a single error of one type. In fact, this code is of distance $2$, e.g. the $X_{2^{r-1}-1}Z_{2^{r-1}}$ operator cannot be detected.
@@ III Quantum dual extended Hamming code
$g_{0}=$ $\bigotimes_{j=2^{r-1}}^{2^{r}-1}\;X_{j}\,Z_{2^{r}-1-j}\,X_{j+2^{r}},$ (7a)
$g_{i}=$ $\bigotimes_{j=0}^{2^{i}-1}\bigotimes_{k=0}^{2^{r-i}-1}\;X_{j(2^{r-i})+k}\,Z_{2^{r}-1-(j(2^{r-i})+k)}\,X_{j(2^{r-i})+k+2^{r}},$
@@
logical qubits. The aforementioned weight-2 operator $X_{2^{r-1}-1}Z_{2^{r-1}}$ is now becomes $X_{2^{r-1}-1}Z_{2^{r-1}}X_{2^{r-1}-1+2^{r}}$, and therefore this code is of distance $3$ rather than the desired distance $4$ of the classical extended Hamming code.
@@ IV Conclusion
In this work, we considered scenarios where only a limited number of high-quality qubits are available, such as in quantum memories and communication systems subject to unavoidable CC errors and minor stochastic Pauli errors. To address these challenges, we introduced a new family of CE stabilizer codes with parameters $2^{r+1},2^{r}-(r+1),3]]$, designed to mitigate both error types while optimizing code rate. The smallest member of this family, the $[[8,1,3]]$ CE code, reduces the required qubit count by half compared to the previously known CE code and improves the error threshold by a factor of four.